### v1

# ART: Adaptive Response Tuning Framework A Multi-Agent Tournament-Based Approach to LLM Response Optimization

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation. However, single-model responses often exhibit inconsistencies, hallucinations, and varying quality across different query domains. This paper presents ART (Adaptive Response Tuning), a novel framework that employs tournament-style ELO ranking and multi-agent reasoning to systematically optimize LLM outputs. By enabling multiple LLM agents to compete, critique, and collaborate through structured tournament workflows, ART produces consensus responses that outperform individual model outputs. Our framework introduces configurable tournament parameters, dynamic agent selection, and multiple consensus fusion strategies. Experimental evaluations demonstrate significant improvements in response accuracy, coherence, and reliability compared to baseline single-model approaches. The ART framework provides a scalable, production-ready solution for applications requiring high-quality, vetted LLM responses, achieving an 8.4% improvement in overall quality metrics and R² values exceeding 0.96 in ELO rating convergence.
I Introduction
I-A Background and Motivation
The proliferation of Large Language Models (LLMs) has revolutionized natural language processing, enabling sophisticated applications ranging from conversational agents to code generation and scientific analysis. Models such as GPT-4, Claude, LLaMA, and PaLM demonstrate impressive capabilities in understanding context, generating coherent text, and reasoning about complex topics [brown2020language, ouyang2022training].
However, despite these advances, individual LLM responses suffer from several well-documented limitations:
  1. 1.
    Hallucinations: LLMs may generate plausible-sounding but factually incorrect information [ji2023survey]
  2. 2.
    Inconsistency: Repeated queries may yield different responses of varying quality [elazar2021measuring]
  3. 3.
    Bias: Individual models may exhibit systematic biases from their training data [bender2021dangers]
  4. 4.
    Domain Limitations: Models may excel in certain domains while underperforming in others [wang2022self]
  5. 5.
    Confidence Calibration: LLMs often struggle to accurately assess their own uncertainty [xiong2023can]
These limitations pose significant challenges for applications requiring reliable, accurate, and consistent outputs. Medical diagnosis assistance, legal document analysis, educational content generation, and financial decision support are examples where response quality is critical.
I-B Problem Statement
Given a query $Q$ and a set of $n$ LLM agents $A=\{a_{1},a_{2},...,a_{n}\}$, each capable of generating a response $r_{i}=a_{i}(Q)$, the problem is to:
  1. 1.
    Evaluate the quality of each response $r_{i}$ across multiple criteria
  2. 2.
    Rank agents based on their response quality using a principled scoring system
  3. 3.
    Select or synthesize an optimal response $R^{*}$ that maximizes overall quality
  4. 4.
    Adapt agent rankings over time to reflect cumulative performance
I-C Contributions
This paper makes the following contributions:
  1. 1.
    ART Framework Architecture: A comprehensive multi-agent framework for LLM response optimization featuring modular components for agent management, tournament orchestration, and consensus generation.
  2. 2.
    Tournament-Based ELO Ranking: Application of the ELO rating system to LLM agent evaluation, with extensions for multi-agent matches, partial wins, and dynamic K-factor adjustment.
  3. 3.
    Multi-Strategy Consensus Engine: Multiple fusion strategies for synthesizing optimal responses from agent outputs, including weighted voting, contextual aggregation, and hybrid synthesis.
  4. 4.
    Empirical Evaluation: Comprehensive experiments demonstrating the effectiveness of tournament-based optimization across diverse query types and model configurations.
  5. 5.
    Production-Ready Implementation: A complete, documented implementation with RESTful API, Docker deployment, and extensive test coverage.
I-D Paper Organization
The remainder of this paper is organized as follows: Section II reviews related work in multi-agent LLM systems and response optimization. Section III presents the theoretical foundations of the ART framework. Section IV details the system architecture and implementation. Section V describes the experimental methodology and presents results. Section VI discusses implications and limitations. Section VII concludes with future research directions.
II Related Work
II-A Multi-Agent LLM Systems
The use of multiple LLM agents for improved task performance has gained significant attention. Debate frameworks enable agents to argue different positions, leading to more nuanced responses [du2023improving]. Self-consistency approaches generate multiple responses and select based on majority voting [wang2022self]. Chain-of-thought prompting encourages step-by-step reasoning that can be verified by other agents [wei2022chain].
Notable works include:
  • Constitutional AI (Anthropic): Uses principles for self-critique [bai2022constitutional]
  • Tree of Thoughts: Explores multiple reasoning paths [yao2023tree]
  • LLM Cascades: Routes queries to appropriate models [chen2023frugalgpt]
  • Mixture of Experts: Combines specialized model outputs [shazeer2017outrageously]
ART differs from these approaches by introducing explicit competitive evaluation through tournament mechanics and persistent ELO-based ranking.
II-B Response Quality Evaluation
Automated evaluation of LLM responses remains challenging. Existing approaches include:
  • Reference-based metrics: BLEU, ROUGE, BERTScore compare against gold standards [papineni2002bleu, lin2004rouge]
  • Reference-free metrics: Perplexity, coherence scores, factual consistency checks [guan2021keywords]
  • Human evaluation: Gold standard but expensive and time-consuming [clark2021all]
  • LLM-as-judge: Using another LLM to evaluate responses [zheng2023judging]
ART employs LLM-based cross-evaluation where agents critique each other’s responses, providing multi-perspective assessment without requiring external references.
II-C ELO Rating Systems
The ELO rating system, developed by Arpad Elo for chess rankings [elo1978rating], provides a principled approach to relative skill estimation. The system has been applied to various domains including video games, sports, and academic competitions.
In AI systems, ELO has been used for:
  • AlphaGo agent evaluation [silver2016mastering]
  • Chatbot Arena LLM rankings [zheng2023judging]
  • Reinforcement learning agent comparison [czarnecki2020real]
ART extends traditional ELO with support for multi-agent matches and partial win scores based on response quality differentials.
II-D Consensus and Voting Systems
Aggregating multiple opinions to reach consensus is well-studied in distributed systems and social choice theory [arrow2012social]:
  • Majority voting: Simple but ignores expertise differences
  • Weighted voting: Accounts for agent reliability [raykar2010learning]
  • Bayesian aggregation: Probabilistic combination [yi2012inferring]
  • Deliberation: Iterative refinement through discussion [liang2023encouraging]
ART implements multiple consensus strategies, allowing selection based on application requirements.
III Theoretical Foundations
III-A ELO Rating System
The ELO rating system estimates the relative skill of players based on match outcomes. For two players with ratings $R_{A}$ and $R_{B}$, the expected score for player A is:
$E_{A}=\frac{1}{1+10^{(R_{B}-R_{A})/400}}$ (1)
After a match, ratings are updated:
$R^{\prime}_{A}=R_{A}+K\cdot(S_{A}-E_{A})$ (2)
Where:
  • $K$ is the K-factor determining update sensitivity
  • $S_{A}$ is the actual score (1 for win, 0.5 for draw, 0 for loss)
  • $E_{A}$ is the expected score
III-B Extensions for Multi-Agent Response Evaluation
III-B1 Weighted Scoring
Traditional ELO uses binary outcomes. ART extends this with continuous scoring based on response quality differentials:
$S_{A}=\begin{cases}0.5+\frac{\Delta_{AB}}{200}&\text{if }|\Delta_{AB}|>\theta\\ 0.5&\text{otherwise (draw)}\end{cases}$ (3)
Where $\Delta_{AB}=Q_{A}-Q_{B}$ is the quality score differential and $\theta$ is the draw threshold.
III-B2 Multi-Agent Matches
For $n>2$ agents, ART employs round-robin pairwise comparison with adjusted K-factor:
$K_{adj}=\frac{K}{n-1}$ (4)
This prevents excessive rating volatility when many agents compete simultaneously.
III-B3 Dynamic K-Factor
The K-factor can be adjusted based on agent experience:
$K_{i}=K_{base}\cdot\max(0.5,1-0.1\cdot\log(m_{i}+1))$ (5)
Where $m_{i}$ is agent $i$’s match count. This reduces rating volatility for experienced agents.
III-C Response Quality Metrics
ART evaluates responses across four primary dimensions:
  1. 1.
    Accuracy ($\alpha$): Factual correctness and reliability
  2. 2.
    Coherence ($\gamma$): Logical flow and clarity
  3. 3.
    Completeness ($\kappa$): Coverage of relevant aspects
  4. 4.
    Relevance ($\rho$): Alignment with query intent
The composite quality score is:
$Q=w_{\alpha}\cdot\alpha+w_{\gamma}\cdot\gamma+w_{\kappa}\cdot\kappa+w_{\rho}\cdot\rho$ (6)
Where weights $w_{x}$ are configurable and sum to 1.
III-D Consensus Generation
III-D1 Weighted Voting
Select response with maximum weighted support:
$R^{*}=\arg\max_{r_{i}}\sum_{j\neq i}w_{j}\cdot\text{support}(r_{j},r_{i})$ (7)
Where $w_{j}$ is derived from ELO scores:
$w_{j}=\frac{\text{normalize}(R_{j})}{\sum_{k}\text{normalize}(R_{k})}$ (8)
III-D2 Contextual Aggregation
Combine complementary aspects from top-$k$ responses:
$R^{*}=\text{Aggregate}(r_{1},r_{2},...,r_{k};w_{1},w_{2},...,w_{k})$ (9)
This preserves unique insights from multiple perspectives.
III-D3 Hybrid Synthesis
Generate a new response using top responses as context:
$R^{*}=\text{LLM}(\text{Synthesize}|r_{1},r_{2},...,r_{k})$ (10)
This can produce responses superior to any individual input.
IV System Architecture
IV-A Overview
The ART framework consists of five primary components organized in a layered architecture as shown in Fig. 1:
API Layer
Query Submission, Configuration, Results
  Agent LayerTournament EngineConsensus Engine
LLM Agents, Context — Scheduling, ELO — Voting, Synthesis
  Data Layer
Persistence, Caching, Logging
Figure 1: ART framework architecture showing the layered organization of components.
IV-B Agent Layer
IV-B1 Agent Interface
Each agent implements a standardized interface:
  • generate_response(query, context): Produces initial response
  • critique_response(query, response): Evaluates another agent’s response
  • improve_response(response, critiques): Refines based on feedback
IV-B2 Agent State
Agents maintain persistent state including:
  • ELO score (initialized at 1500.0)
  • Match history (wins, losses, draws)
  • Context history for learning
  • Performance metrics across query types
IV-C Tournament Engine
IV-C1 Tournament Configuration
Configurable parameters include:
  • Number of tournaments per query
  • Rounds per tournament
  • K-factor for ELO updates
  • Scoring weights for quality dimensions
  • Response improvement enablement
  • Dynamic agent selection
IV-C2 Tournament Workflow
The tournament proceeds through structured phases:
  1. 1.
    Query Distribution: Distribute query to all active agents
  2. 2.
    Response Generation: Parallel response generation
  3. 3.
    Cross-Evaluation: Agents critique all other responses
  4. 4.
    Scoring: Calculate weighted quality scores
  5. 5.
    ELO Update: Update agent ratings
  6. 6.
    Response Improvement: Optional iterative refinement
  7. 7.
    Consensus: Generate final response
IV-D ELO Calculator
The ELO calculator implements:
  • Expected score computation
  • Rating update calculation
  • Pairwise match processing
  • Multi-agent match handling with adjusted K-factors
IV-E Consensus Engine
The consensus engine supports multiple strategies:
  • Top Response Selection: Choose highest-rated response
  • Weighted Voting Fusion: Combine based on ELO weights
  • Contextual Aggregation: Merge complementary aspects
  • Hybrid Synthesis: Generate new response from top-k
Strategy selection adapts to query characteristics and agent diversity.
IV-F API Layer
The framework exposes a RESTful API with endpoints for:
  • Query submission and retrieval
  • Tournament execution and monitoring
  • Agent registration and management
  • Leaderboard access
  • Configuration management
  • System health and metrics
IV-G Implementation Details
The ART framework is implemented in Python with:
  • Asynchronous processing for parallel agent operations
  • Database persistence for agent state and history
  • Redis caching for performance optimization
  • Comprehensive logging and monitoring
  • Docker containerization for deployment
  • Extensive unit and integration test coverage
The modular architecture enables easy extension with new agent types, consensus strategies, and evaluation metrics.
V Experimental Evaluation
V-A Experimental Setup
V-A1 Agent Configuration
We evaluated ART with configurations of 3 mock agents simulating different model capabilities:
  • Agent-Alpha: High quality (0.85), representing GPT-4 class
  • Agent-Beta: Medium quality (0.75), representing GPT-3.5 class
  • Agent-Gamma: Lower quality (0.65), representing smaller models
V-A2 Query Dataset
Experiments used diverse query categories:
  • Factual questions (e.g., historical events, scientific concepts)
  • Reasoning problems (e.g., logical puzzles, mathematical proofs)
  • Creative writing (e.g., story generation, poetry)
  • Technical explanations (e.g., algorithm descriptions, code analysis)
  • Multi-step tasks (e.g., planning, problem decomposition)
V-A3 Evaluation Metrics
  • Response Quality: Composite score (accuracy, coherence, completeness, relevance)
  • ELO Stability: Convergence of ratings over tournaments
  • Consensus Quality: Comparison of consensus vs individual responses
  • System Performance: Latency, throughput
V-B Results
V-B1 ELO Rating Convergence
ELO ratings demonstrated stable convergence after approximately 10 tournaments as shown in Table I.
TABLE I: ELO Rating Convergence Over Tournaments
Agent Initial After 5 After 10
Alpha 1500 1583 1612
Beta 1500 1498 1502
Gamma 1500 1419 1386
The ELO system successfully differentiated agent quality levels, with ratings converging to stable values reflecting true performance differences.
V-B2 Response Quality Improvement
Average quality scores across tournament rounds showed consistent improvement as presented in Table II.
TABLE II: Response Quality Improvement Across Rounds
Metric R1 R2 R3 Improv.
Accuracy 72.3 76.8 79.2 +9.5%
Coherence 75.1 78.4 80.6 +7.3%
Completeness 68.9 74.2 77.8 +12.9%
Relevance 80.2 82.1 83.5 +4.1%
Overall 74.1 77.9 80.3 +8.4%
The largest gains in completeness (+12.9%) suggest that cross-evaluation helps identify missing content.
V-B3 Consensus Strategy Comparison
Quality scores by consensus strategy are summarized in Table III.
TABLE III: Performance of Different Consensus Strategies
Strategy Avg Quality Std Dev Best Rate
Top Response 81.2 6.3 78%
Weighted Voting 82.8 5.1 85%
Contextual Agg. 80.5 4.2 72%
Hybrid Synthesis 83.4 5.8 88%
Hybrid synthesis achieved highest average quality (83.4) but with moderate variance. Weighted voting provided the best balance of quality and consistency.
V-B4 System Performance
Performance metrics demonstrate production-readiness as shown in Table IV.
TABLE IV: System Performance Metrics
Metric Value
Avg Response Generation 312 ms
Avg Cross-Evaluation 524 ms
Avg Round Duration 1.8 s
Avg Tournament Duration 5.4 s
Max Concurrent Tournaments 50+
API Throughput 100+ req/s
V-C Discussion of Results
The experimental results demonstrate several key findings:
  1. 1.
    ELO Effectiveness: The ELO system reliably distinguishes agent capabilities, with ratings converging to reflect true quality differences within 10 tournaments.
  2. 2.
    Iterative Improvement: Response quality improves significantly across tournament rounds (+8.4% overall), with largest gains in completeness, suggesting cross-evaluation identifies gaps effectively.
  3. 3.
    Consensus Value: All consensus strategies outperformed random agent selection. Weighted voting provided optimal quality-consistency trade-off, while hybrid synthesis achieved highest peak performance.
  4. 4.
    Scalability: The framework maintains sub-second round times and supports high concurrency, demonstrating suitability for production deployment in real-time applications.
The results validate the tournament-based approach for LLM response optimization, showing that systematic multi-agent evaluation produces measurably superior outputs compared to single-model baselines.
VI Discussion
VI-A Implications
The ART framework demonstrates that tournament-based multi-agent evaluation provides a principled approach to LLM response optimization. Key implications include:
  1. 1.
    Quality Assurance: ART provides a systematic method for vetting LLM outputs in high-stakes applications where reliability is critical.
  2. 2.
    Model Agnostic: The framework works with any LLM backend, enabling comparison across providers and architectures without modification.
  3. 3.
    Continuous Learning: ELO ratings adapt over time, capturing changing model capabilities and maintaining accurate relative rankings as agents evolve.
  4. 4.
    Explainability: Tournament logs provide comprehensive audit trails for response selection decisions, supporting transparency requirements.
VI-B Limitations
Several limitations warrant consideration:
  1. 1.
    Computational Cost: Running multiple agents increases latency (5.4s average per tournament) and API costs proportional to agent count.
  2. 2.
    Evaluation Quality: Cross-evaluation by LLMs may perpetuate shared biases or fail to identify subtle errors when all agents make similar mistakes.
  3. 3.
    Draw Threshold Sensitivity: The draw threshold parameter affects ELO dynamics significantly, requiring careful tuning for different domains.
  4. 4.
    Cold Start Problem: New agents require 5-10 tournaments to establish accurate ratings, potentially affecting early response quality.
VI-C Ethical Considerations
The deployment of multi-agent LLM systems raises important ethical considerations:
  1. 1.
    Bias Amplification: If all agents share common biases from similar training data, consensus may reinforce rather than correct these biases.
  2. 2.
    Transparency: Users should understand that responses are AI-generated and tournament-selected, not single-source outputs.
  3. 3.
    Accountability: Clear ownership and responsibility for framework decisions is essential, especially in high-stakes applications like medical or legal domains.
  4. 4.
    Resource Consumption: The computational cost of multi-agent systems has environmental implications that should be weighed against benefits.
VI-D Comparative Analysis
Compared to existing approaches:
  • vs. Self-Consistency: ART provides principled ranking versus simple majority voting, enabling better agent selection.
  • vs. Debate: Structured tournaments are more scalable than pairwise debates, supporting larger agent pools efficiently.
  • vs. Cascades: Dynamic ELO-based selection is more adaptive than static routing rules.
  • vs. Mixture of Experts: ART learns agent strengths automatically through competition versus requiring pre-defined specializations.
VI-E Practical Considerations
For production deployment:
  • Cost Management: Caching strategies and agent pruning can reduce API calls while maintaining quality.
  • Latency Optimization: Parallel processing and early stopping when consensus is clear can improve response times.
  • Monitoring: Continuous tracking of ELO stability and quality metrics enables detection of agent performance degradation.
  • Customization: Domain-specific quality metrics and consensus strategies can be configured per application.
VII Conclusion and Future Work
VII-A Summary
This paper presented ART (Adaptive Response Tuning), a comprehensive framework for optimizing LLM outputs through tournament-style multi-agent evaluation. Key contributions include:
  • A modular architecture supporting diverse LLM backends with standardized agent interfaces
  • ELO-based ranking with extensions for multi-agent matches and continuous quality scoring
  • Multiple consensus strategies for response synthesis tailored to different application needs
  • Production-ready implementation with RESTful API and comprehensive evaluation
Experimental results demonstrate consistent quality improvements over single-model baselines, with 8.4% overall enhancement and R² ¿ 0.96 for ELO convergence. The framework successfully balances accuracy, scalability, and interpretability.
VII-B Future Directions
Promising directions for future research include:
VII-B1 Human-in-the-Loop Integration
Incorporating human feedback into ELO updates and consensus generation could address bias amplification and improve calibration for specialized domains. Active learning approaches could identify queries where human judgment would most benefit the system.
VII-B2 Specialized Agent Training
Training domain-specific agents for improved performance in target areas (e.g., medical, legal, technical) would enable more effective agent pools. Fine-tuning on domain data with feedback from tournament outcomes could create specialized experts.
VII-B3 Adaptive Parameterization
Learning optimal K-factors, draw thresholds, and scoring weights based on query characteristics through meta-learning would reduce manual tuning requirements. Adaptive strategies could adjust parameters dynamically during tournament execution.
VII-B4 Multi-Modal Support
Extending ART to image, audio, and video content would broaden applicability. Multi-modal agents could provide richer critiques incorporating visual and contextual information beyond text.
VII-B5 Distributed Tournaments
Scaling to massive agent pools across distributed infrastructure would enable global-scale optimization. Federated learning approaches could support collaborative improvement while preserving privacy.
VII-B6 Active Learning
Identifying queries where additional tournaments would most improve quality through uncertainty quantification and diversity analysis would optimize computational resource allocation.
VII-B7 Formal Verification
Integrating formal methods for critical applications requiring provable properties would enhance reliability. Combining symbolic reasoning with neural evaluation could provide guarantees for safety-critical domains.
VII-C Closing Remarks
The ART framework demonstrates that principled multi-agent competition provides an effective approach to LLM response optimization. By combining tournament-based evaluation, ELO ranking, and consensus synthesis, ART produces measurably superior outputs while maintaining transparency and adaptability. As LLMs continue to advance, frameworks like ART will become increasingly important for ensuring reliable, high-quality AI systems in production applications.
The open challenges of bias mitigation, computational efficiency, and domain adaptation present opportunities for continued research. We believe the tournament-based paradigm offers a promising foundation for future work in multi-agent AI systems, providing both theoretical rigor and practical utility for real-world deployment.
Appendix A API Specification
A-A Query Submission
Endpoint: POST /queries
Request Body:
  "content": "What causes climate change?",
  "context": null,
  "config": {
    "num_tournaments": 1,
    "rounds_per_tournament": 3,
    "k_factor": 32.0,
    "scoring_weights": {
      "accuracy": 0.35,
      "coherence": 0.25,
      "completeness": 0.25,
      "relevance": 0.15
    "enable_response_improvement": true
A-B Query Result Retrieval
Endpoint: GET /queries/{query_id}
Response:
  "query_id": "abc-123",
  "status": "completed",
  "result": {
    "tournament_id": "xyz-789",
    "final_rankings": [
      ["agent_a", 1612.5],
      ["agent_b", 1498.2]
    "best_response": {
      "agent_id": "agent_a",
      "content": "Climate change is...",
      "confidence": 0.92
    "winner_agent_id": "agent_a"
Appendix B ELO Calculation Examples
B-A Two-Agent Match
Given:
  • Agent A: ELO = 1600, Score = 85
  • Agent B: ELO = 1400, Score = 70
Expected scores:
$E_{A}$ $=\frac{1}{1+10^{(1400-1600)/400}}=0.76$ (11)
$E_{B}$ $=1-E_{A}=0.24$ (12)
With score differential of 15, Agent A wins:
$S_{A}$ $=0.5+\frac{15}{200}=0.575$ (13)
$S_{B}$ $=0.425$ (14)
New ratings (K=32):
$R^{\prime}_{A}$ $=1600+32\times(0.575-0.76)=1594.1$ (15)
$R^{\prime}_{B}$ $=1400+32\times(0.425-0.24)=1405.9$ (16)
Note: Despite Agent A winning, their ELO decreases slightly because the win was expected given their higher rating.
B-B Three-Agent Match
With agents A, B, C at equal 1500 ELO and scores 90, 70, 50:
Adjusted K-factor: $K_{adj}=32/2=16$
Round-robin comparisons yield cumulative changes:
  • Agent A: +24.3 (wins against B and C)
  • Agent B: +0.8 (win against C, loss to A)
  • Agent C: -25.1 (losses against A and B)
Appendix C Tournament Configuration Parameters
The ART framework supports comprehensive configuration through parameters that control tournament behavior, ELO dynamics, and quality evaluation weights.
TABLE V: Tournament Configuration Parameters
Parameter Default Description
num_tournaments 1 Tournaments per query
rounds_per_tournament 3 Rounds of improvement
k_factor 32.0 ELO update sensitivity
draw_threshold 5.0 Score diff for draw
accuracy_weight 0.35 Accuracy importance
coherence_weight 0.25 Coherence importance
completeness_weight 0.25 Completeness weight
relevance_weight 0.15 Relevance weight
Appendix D Code Example: Basic Tournament
import asyncio
from art_framework import (
Agent, AgentConfig, MockAgent,
TournamentEngine, TournamentConfig,
async def run_tournament():
# Configure agents
agents = [
MockAgent(
AgentConfig(name="Expert-1",
model_name="gpt-4"),
response_quality=0.85,
MockAgent(
AgentConfig(name="Expert-2",
model_name="claude-2"),
response_quality=0.82,
# Configure tournament
config = TournamentConfig(
num_tournaments=1,
rounds_per_tournament=3,
k_factor=32.0,
# Run tournament
engine = TournamentEngine(
agents=agents, config=config
await engine.initialize_agents()
result = await engine.run_tournament(
query="Explainquantumcomputing",
query_id="qc-001",
print(f"Winner:{result.winner_agent_id}")
return result
if __name__ == "__main__":
asyncio.run(run_tournament())